Direct Materials Purchases Budget Anticipated sales for Safety Grip Company were 40,000 passenger car tires and...

  1. Direct Materials Purchases Budget

    Anticipated sales for Safety Grip Company were 40,000 passenger car tires and 12,000 truck tires. Rubber and steel belts are used in producing passenger car and truck tires according to the following table:

    Passenger Car Truck
    Rubber 31 lbs. per unit 72 lbs. per unit
    Steel belts 7 lbs. per unit 18 lbs. per unit

    The purchase prices of rubber and steel are $3.1 and $4.1 per pound, respectively. The desired ending inventories of rubber and steel belts are 38,000 and 8,000 pounds, respectively. The estimated beginning inventories for rubber and steel belts are 44,000 and 6,000 pounds, respectively.

    Prepare a direct materials purchases budget for Safety Grip Company for the year ended December 31, 20Y9.

    Safety Grip Company
    Direct Materials Purchases Budget
    For the Year Ending December 31, 20Y9
    Rubber Steel Belts Total
    Pounds required for production:
    Passenger tires lbs. lbs.
    Truck tires
    • Desired inventory, December 1, 20Y9
    • Desired inventory, December 31, 20Y9
    • Direct labor hour
    • Estimated inventory, December 1, 20Y9
    • Estimated inventory, December 31, 20Y9
    Total pounds available lbs. lbs.
    • Depreciation
    • Desired inventory, January 1, 20Y9
    • Desired inventory, January 31, 20Y9
    • Estimated inventory, January 1, 20Y9
    • Estimated inventory, January 31, 20Y9
    Total units purchased lbs. lbs.
    Unit price x $ x $
    Total direct materials to be purchased $ $ $

    Feedback

    Remember to take into account materials required for production, desired ending materials inventory and estimated beginning materials inventory when calculating direct materials to be purchased.

    Learning Objective 4.

Loading item

There was an error loading this item. If this continues to occur, please contact Technical Support.

Check My Work

  • Previous
  • Next
  • 100% Correct
  • Partially Correct
  • Incorrect
  • Needs Instructor Grading

Basic Calculatorclose

0

UseEntBSBSpCEHomCEnd

789+

456-

123*

0.=/
Answer #1
Safety Grip Company
Direct Materials Purchases Budget
For the Year Ending December 31, 20Y9
Rubber Steel Belts Total
Pounds required for production:
Passenger tires 1240000 280000
Truck tires 864000 216000
Desired inventory, December 31, 20Y9 38000 8000
Total pounds available 2142000 504000
Estimated inventory, January 1, 20Y9 44000 6000
Total units purchased 2098000 498000
Unit price 3.1 4.1
Total direct materials to be purchased 6503800 2041800 8545600
Estimated inventory, January 1, 20Y9 can alternatively be written with a minus sign
Workings:
Rubber Steel Belts
Pounds required for production:
Passenger tires =40000*31 =40000*7
Truck tires =12000*72 =12000*18